# Rapid formation of a very massive star $>\!\!50000\,\mathrm{M}_{\odot}$ and subsequently an IMBH from runaway collisions

### v1

@@ 3.2 Initial Conditions for MOCCA and Code Modifications
Similar to the direct N-body runs, the input parameters that govern stellar and binary evolution prescriptions in MOCCA were set to level C as described in Kamlah et al. (2022a) and Appendix A. For these runs, the overall time step in MOCCA was reduced to 0.02 Myr. Given the short central relaxation time of this dense model, the smaller time step allowed us to better resolve the evolution and growth of the VMS. Additionally, it enabled more accurate capture of mass segregation and the effects of mass loss on the cluster’s evolution. In the MOCCA run, three-body binary formation (3BBF) was specifically disabled for the VMS only when its mass exceeded $500\ \leavevmode\nobreak\ \rm M_{\odot}$ to allow for a more accurate comparison with N-body results. This modification was necessary because, in MOCCA, a star in a binary is not allowed to undergo two-body hyperbolic collisions. If the VMS is in a binary, it can only grow through collisions or mergers during binary-single and binary-binary encounters. With 3BBF enabled for the VMS, it frequently formed a binary and remained bound for extended periods, preventing it from growing through two-body collisions. As a result, in the simulation with 3BBF enabled, the VMS reached a mass of only 35,000 $\rm M_{\odot}$ within 4.5 Myr, significantly limiting its growth compared to the direct N-body simulation. This restriction is a limitation of MOCCA that does not exist in NBODY6++GPU, where two-body collisions can still occur even if the VMS is part of a binary system. By disabling 3BBF for the VMS in MOCCA, we allowed it to undergo two-body collisions more freely, leading to a growth rate much closer to that observed in NBODY6++GPU. This adjustment was crucial for achieving a more consistent comparison between the two simulation methods. Recent work by Atallah et al. (2024) on 3BBF showed that in interactions with a large mass ratio, 3BBF favors pairing the two least massive bodies. Given these findings, disabling 3BBF for the VMS in MOCCA may not be an unreasonable approach, as it prevents the VMS from being preferentially locked into a binary with a much lower-mass companion, thereby allowing for more efficient growth through collisions.
@@ 4.1 Dynamical star cluster evolution and the evolution of a VMS and formation of an IMBH
We dynamically evolve our system for 555\,5Myr, when already a VMS had formed and evolved into an IMBH. However, relaxation processes are still relevant, since mass segregation takes place much faster than the global half-mass relaxation time given above, with a difference of three order of magnitude (Khalisi et al., 2007).
@@
Figure 2: The figure shows the average mass in each shown Lagrangian shells $\langle M\rangle_{\mathrm{Lagr}}[\mathrm{M}_{\odot}]$ (Top) and, the Lagrangian radii $R_{\mathrm{Lagr}}\leavevmode\nobreak\ [\mathrm{pc}]$ (Bottom), both normalised by the cluster mass at the beginning of the simulation. We show both N-body and MOCCA results for the Lagrangian radii. In both plots, we show the shell representing 1 %, 10 %, 30%, 50%, 70% and 90 %. The differences between the Lagrangian radii in the N-body and MOCCA simulations arise from how escaping stars are treated. In the N-body simulations, stars with positive energy are not immediately removed from the system; instead, they remain until they cross a boundary set at ten times the half-mass radius. These stars are still included in the calculation of Lagrangian radii, which affects the results compared to MOCCA.
@@
Figure 2 shows the average mass of the cluster inside the Lagrangian radii (top) and the Lagrangian radii (bottom), which shows the different evolution of shells from the inner to the outer regions. In our model, the star cluster evolves rapidly in the innermost regions ($1\%$), where strong mass segregation triggers an early core collapse. The $10\%$ Lagragian radii start to decrease after 333\,3Myr of evolution. The $30\%$, $50\%$, $70\%$ and $90\%$ Lagrangian radii show the smooth expansion of the cluster; Until around 4.694.694.69\,4.69Myr (solid lines) and 4.484.484.48\,4.48Myr (dashed lines) when the IMBH is formed, in N-body and MOCCA simulations, respectively. For the innermost mass shell (1%) we can follow the formation of the VMS, as its mass is getting larger than the corresponding Lagrangian shell’s mass. Notably, the VMS reaches a mass of $M_{\rm VMS}\geq 1000\leavevmode\nobreak\ \rm M_{\odot}$ at time 0.180.180.18\,0.18Myr (N-body simulation) and 0.220.220.22\,0.22Myr (MOCCA simulation).
@@
Figure 3 shows the cumulative mass of the escapers during the simulation. In the first Myrs, the cumulative escaped mass is approximately 4 0005 $4\,000-5\,000\leavevmode\nobreak\ \rm M_{\odot}$, increasing to around 20 $20\,000\leavevmode\nobreak\ \rm M_{\odot}$ after 555\,5Myr. N-body and MOCCA simulation show reasonable agreement. The figure also presents the total number of collisions, which amounts to 22 0102201022\,01022 010. Of these, $91.6\%$ involves the VMS/IMBH, these collisions can be divided into binary collisions ($45.1\%$) and hyperbolic collisions ($46.5\%$). Binary collisions are more frequent when the VMS mass exceeds 1 $1\,000\leavevmode\nobreak\ \rm M_{\odot}$ at time 0.180.180.18\,0.18Myr. In contrast, hyperbolic collisions become more frequent as the VMS grows in size, reaching a radius larger than 1 $1\,000\leavevmode\nobreak\ \rm R_{\odot}$ at time 3.363.363.36\,3.36Myr. For the MOCCA simulation, we do not distinguish between binary and hyperbolic collisions, as the code does not explicitly track stellar orbits. Some collisions may involve stars that were gravitationally bound to the VMS, but confirming this would require reconstructing stellar orbits from snapshots using their instantaneous positions, velocities, and the local gravitational potential, which is challenging due to the limited temporal resolution. Consequently, we report only the total number of collisions ($16103$ in 5 Myr), of which $65.6\%$ involve the VMS/IMBH. Despite this limitation, the results from both codes are in good agreement.
@@
Figure 4 compares the NBODY6++GPU and MOCCA simulations, showing the time evolution of the masses of escapers from the cluster (top panels) and secondary stars involved in collisions with the VMS/IMBH (bottom panels). The left panels display the N-body results, and the right panels show the MOCCA results. The color scale represents the local point density of events, with yellow regions indicating higher concentrations of recorded data points. In both simulations, low-mass stars (typically $\lesssim 1\leavevmode\nobreak\ \rm M_{\odot}$) dominate the population of escapers. The N-body simulation shows two prominent epochs of escaper activity: an extended early phase over the first $\sim 2\,$Myr, followed by a shorter but more intense burst near the time of IMBH formation at 4.694.694.69\,4.69Myr. In contrast, the MOCCA simulation exhibits a smoother and more gradual decline in escaper activity over time, with a modest peak near $4.48$ Myr corresponding to IMBH formation.
@@
The lower panels show the masses of stars colliding with the VMS/IMBH. MOCCA displays a broader and more uniform spread across the stellar mass range from the start of the simulation, while the N-body simulation shows occasional high-mass collisions (with $m\gtrsim 10\leavevmode\nobreak\ \rm M_{\odot}$), though these are relatively sparse. In both simulations, most collisions involve low-mass stars. However, these frequent low-mass mergers contribute less significantly to the final VMS/IMBH mass (see also Figure 5). The N-body simulation shows a brief dip in the number of secondary stars involved in collisions around 1.681.681.68\,1.68Myr, lasting approximately 0.10.10.1\,0.1Myr. This feature coincides with the formation of a hard binary between the VMS and a main-sequence star (see Figure 7). Subsequent three-body interactions lead to enhanced escape rates. Toward the end of the simulation, the IMBH sequentially forms two hard binaries with $\sim 2\leavevmode\nobreak\ \rm M_{\odot}$ MS stars, which explains the observed drop in further collisions.
@@
Figure  5 shows histograms of the number of collisions (top panels) and the cumulative mass of the secondary star colliding with the VMS/IMBH (bottom panels). The N-body results are shown in the left panels, while the MOCCA results are shown in the right panels, the histogram mass ranges $<0.1$, $0.1-1$, $1-10$, $10-100$ and $>100\leavevmode\nobreak\ \rm M_{\odot}$. The number of collisions for N-body are $2057$, $14525$, $2525$, $923$, $95$, while for MOCCA are $1053$, $7797$, $2273$, $1570$, $74$, for the first two mass ranges, N-body has about twice as many collisions as MOCCA, for the mass range $1-10\leavevmode\nobreak\ \rm M_{\odot}$ the number of collisions is about 300 more in N-body, while for the next mass range $10-100\leavevmode\nobreak\ \rm M_{\odot}$ it is higher for MOCCA, about 600 more collisions. Finally, for the last mass range $>100\leavevmode\nobreak\ \rm M_{\odot}$ the number of collisions is slightly large in the N-body simulation. In the case of N-body simulations where it is possible to distinguish between binary (red bars) and hyperbolic (blue bars) collisions, we can observe that hyperbolic collisions contribute slightly more to the VMS formation for the mass range $<0.1$, $0.1-1\leavevmode\nobreak\ \rm M_{\odot}$. The mass range $1-10\leavevmode\nobreak\ \rm M_{\odot}$ shows almost the same number of collisions, however, binary collisions contribute slightly more to the VMS mass, while binary collisions contribute quite a bit more to the VMS formation for the mass range $10-100$, and $>100\leavevmode\nobreak\ \rm M_{\odot}$. Both codes shows the largest number of collisions comes from stars in the mass range $0.1-1\leavevmode\nobreak\ \rm M_{\odot}$. However, the largest mass contribution to the VMS comes from stars in the mass range $10-100\leavevmode\nobreak\ \rm M_{\odot}$, for MOCCA the mass contribution is $4.41\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, while in N-body it is $3.35\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$ ($2.73\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$ from binary collisions (cyan bar) and $6.34\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ from hyperbolic collisions (magenta bar)), the second large mass contribution to the VMS comes from the mass range $>100\leavevmode\nobreak\ \rm M_{\odot}$, in MOCCA with $9.46\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ and in N-body with $1.26\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, with $1.17\times 10^{4}$ and $914.19\leavevmode\nobreak\ \rm M_{\odot}$, from binary and hyperbolic collisions, respectively. For the lower mass ranges $<0.1$, $0.1-1$ and, $1-10\leavevmode\nobreak\ \rm M_{\odot}$ the VMS mass contribution in the same order is $94.49$, $2.59\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ and, $8.18\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ for MOCCA, while for N-body it is $184.01$, $4.69\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ and, $6.80\times 10^{3}\leavevmode\nobreak\ \rm M_{\odot}$, for this mass ranges the VMS mass contribution from binary and hyperbolic collisions is quite similar.
[…]
@@ 4.2 N-body results: VMS evolution and IMBH formation
In our N-body simulation, several collisions along the simulation to build up a VMS with a mass of $56996.91\leavevmode\nobreak\ \rm M_{\odot}$, eventually collapsing and forming an IMBH with a mass of $51625.99\leavevmode\nobreak\ \rm M_{\odot}$ at time 4.694.694.69\,4.69Myr.
@@
Figure 6 shows the scatter of the position relative to the star cluster’s density center of the VMS (and IMBH thereafter), $\xi_{\mathrm{VMS,\,IMBH}}\leavevmode\nobreak\ \rm[pc]$, and the speed of the VMS (and IMBH, when formed), $\nu_{\mathrm{VMS,\,IMBH}}\leavevmode\nobreak\ [\mathrm{km\,s}^{-1}]$. Initially, the star with a mass of $130.58\leavevmode\nobreak\ \rm M_{\odot}$, which later becomes the VMS, is located at $0.088\rm pc$ from the center, sinking with a velocity of $32.9\leavevmode\nobreak\ \rm{km\,s}^{-1}$. After around 100 years, it moves closer to the center at $0.061\rm pc$, while its velocity increases nearly threefold to $99.7\leavevmode\nobreak\ \rm{km\,s}^{-1}$. A few hundreds years later, it reaches a velocity of $165.1\leavevmode\nobreak\ \rm{km\,s}^{-1}$ at a position of $0.018\rm pc$. This rapid infall toward the center is driven by the high cluster density, however, the VMS position and velocity oscillations are consistent with the orbital time. After two collisions with MS stars (i.e., stars with a fraction of a solar mass), the first significant collision occurs with a star of mass $54.89\leavevmode\nobreak\ \rm M_{\odot}$ following a binary interaction. At this point, the star is at $0.006\rm pc$. As shown in the figure, the motion of the VMS, in terms of position and velocity, first aligns with the expectations of mass segregation (steady decrease in distance to the center and a velocity profile that initially rises before declining) and later transitions to Brownian motion. At 1.681.681.68\,1.68Myr a hard binary system forms between the VMS and an MS star. Figure 7 shows the semi-mayor axis and eccentricity of the hard binary formed between the VMS with a mass of 28 $28\,891\leavevmode\nobreak\ \rm M_{\odot}$ and a secondary star with a mass of $75\leavevmode\nobreak\ \rm M_{\odot}$. Both stars are tightly bound due to their small average distance ($<3$ AU), they have an elliptical orbit, this dynamical interaction lasted about 0.10.10.1\,0.1Myr. The binary system have a higher binding energy than the kinetic energy of the surrounding stars, which eventually gain kinetic energy, increasing their velocity. The latter reduces the probability of low-velocity stars that could collide with the VMS. Throughout the simulation, the velocity and position of the VMS remain nearly constant, fluctuating around an average value of the order of a $10^{-3}\rm pc$, while the velocity remains in moderate motion, at a few kilometers per second.
@@
Figure 7: The Figure shows the temporal evolution of a hard binary system during between 1.68 and 1.80 Myr. This binary has the primary star as the VMS, with a mass of $28891\leavevmode\nobreak\ \rm M_{\odot}$ and a secondary star with a mass of $75\leavevmode\nobreak\ \rm M_{\odot}$. This interaction lasts approximately $0.1$ Myr. The top panel shows the semi-major axis, while the bottom panel shows the eccentricity.
@@
As supported by the theoretical expectations discussed in previous sections, the VMS growth in the N-body simulation occurs on a short time-scale. The VMS original progenitor passes from an initial mass of $130.58\leavevmode\nobreak\ \rm M_{\odot}$ to exceed $10^{3}\leavevmode\nobreak\ \rm M_{\odot}$ after only 0.180.180.18\,0.18Myr through 17 collisions, further reading $>5000\leavevmode\nobreak\ \rm M_{\odot}$ in a time 0.290.290.29\,0.29Myr. The VMS continues its growth for 0.480.480.48\,0.48Myr through several collisions until reaching more than $10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, naturally, the size of the VMS also increases having $71.13\rm R_{\odot}$, the continued collisions allow a strong rejuvenation of the star which has an effective age 0.230.230.23\,0.23Myr. Lately, around 1.01.01.0\,1.0Myr the VMS has a mass, effective age, and radius of $1.9\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, 0.540.540.54\,0.54Myr, and $80.75\rm R_{\odot}$, respectively. The evolution of the mass, radius, and age of the VMS is shown in the top, middle, and bottom left panels of Figure 8.
@@
Figure 8: Evolution of the VMS (and IMBH thereafter) over time. Top panel: Mass of the VMS/IMBH, $M_{\mathrm{VMS\,IMBH}}\leavevmode\nobreak\ [\mathrm{M}_{\odot}]$. Middle panel: Stellar radius of the VMS/IMBH, $R_{\mathrm{VMS\,IMBH}}\leavevmode\nobreak\ [\mathrm{R}_{\odot}]$ (logarithmic scale). Bottom panel: Effective age of the VMS during its main-sequence phase, $A_{\mathrm{VMS\,IMBH}}\leavevmode\nobreak\ [\mathrm{Myr}]$.
@@
At 222\,2Myr, the VMS has a mass of $3.1\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, with a radius $146.24\rm R_{\odot}$ and effective age 1.211.211.21\,1.21Myr, the hard binary system formed at 1.681.681.68\,1.68Myr in the N-body simulation prevents the VMS from experiencing further collisions, thus leads to a VMS with about 4 $4\,000\leavevmode\nobreak\ \rm M_{\odot}$ less mass than the VMS obtained in MOCCA simulation, that has a mass of $3.5\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$ at the same simulation time, 111\,1Myr later the N-body VMS reaches a mass, radius, and effective age of $4.3\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$, $507.24\rm R_{\odot}$ and 1.771.771.77\,1.77Myr, respectively. The N-body VMS still have less mass than the MOCCA VMS which have a mass $4.52\times 10^{4}\leavevmode\nobreak\ \rm M_{\odot}$ at 333\,3Myr.

### v2

@@ 3.2 Initial Conditions for MOCCA and Code Modifications
Similar to the direct N-body runs, the input parameters that govern stellar and binary evolution prescriptions in MOCCA were set to level C as described in Kamlah2022a and Appendix A. For these runs, the overall time step in MOCCA was reduced to 0.02 Myr. Given the short central relaxation time of this dense model, the smaller time step allowed us to better resolve the evolution and growth of the VMS. Additionally, it enabled more accurate capture of mass segregation and the effects of mass loss on the cluster’s evolution. In the MOCCA run, three-body binary formation (3BBF) was specifically disabled for the VMS only when its mass exceeded $500\ \penalty 10000\ \rm M_{\odot}$ to allow for a more accurate comparison with N-body results. This modification was necessary because, in MOCCA, a star in a binary is not allowed to undergo two-body hyperbolic collisions. If the VMS is in a binary, it can only grow through collisions or mergers during binary-single and binary-binary encounters. With 3BBF enabled for the VMS, it frequently formed a binary and remained bound for extended periods, preventing it from growing through two-body collisions. As a result, in the simulation with 3BBF enabled, the VMS reached a mass of only 35,000 $\rm M_{\odot}$ within 4.5 Myr, significantly limiting its growth compared to the direct N-body simulation. This restriction is a limitation of MOCCA that does not exist in NBODY6++GPU, where two-body collisions can still occur even if the VMS is part of a binary system. By disabling 3BBF for the VMS in MOCCA, we allowed it to undergo two-body collisions more freely, leading to a growth rate much closer to that observed in NBODY6++GPU. This adjustment was crucial for achieving a more consistent comparison between the two simulation methods. Recent work by Atallah2024 on 3BBF showed that in interactions with a large mass ratio, 3BBF favors pairing the two least massive bodies. Given these findings, disabling 3BBF for the VMS in MOCCA may not be an unreasonable approach, as it prevents the VMS from being preferentially locked into a binary with a much lower-mass companion, thereby allowing for more efficient growth through collisions.
@@ 4.1 Dynamical star cluster evolution and the evolution of a VMS and formation of an IMBH
We dynamically evolve our system for 55\,Myr, when already a VMS had formed and evolved into an IMBH. However, relaxation processes are still relevant, since mass segregation takes place much faster than the global half-mass relaxation time given above, with a difference of three order of magnitude (Khalisi2007).
@@
Figure 2: The figure shows the average mass in each shown Lagrangian shells $\langle M\rangle_{\mathrm{Lagr}}[\mathrm{M}_{\odot}]$ (Top) and, the Lagrangian radii $R_{\mathrm{Lagr}}\penalty 10000\ [\mathrm{pc}]$ (Bottom), both normalised by the cluster mass at the beginning of the simulation. We show both N-body and MOCCA results for the Lagrangian radii. In both plots, we show the shell representing 1 %, 10 %, 30%, 50%, 70% and 90 %. The differences between the Lagrangian radii in the N-body and MOCCA simulations arise from how escaping stars are treated. In the N-body simulations, stars with positive energy are not immediately removed from the system; instead, they remain until they cross a boundary set at ten times the half-mass radius. These stars are still included in the calculation of Lagrangian radii, which affects the results compared to MOCCA.
@@
Figure 2 shows the average mass of the cluster inside the Lagrangian radii (top) and the Lagrangian radii (bottom), which shows the different evolution of shells from the inner to the outer regions. In our model, the star cluster evolves rapidly in the innermost regions ($1\%$), where strong mass segregation triggers an early core collapse. The $10\%$ Lagragian radii start to decrease after 33\,Myr of evolution. The $30\%$, $50\%$, $70\%$ and $90\%$ Lagrangian radii show the smooth expansion of the cluster; Until around 4.694.69\,Myr (solid lines) and 4.484.48\,Myr (dashed lines) when the IMBH is formed, in N-body and MOCCA simulations, respectively. For the innermost mass shell (1%) we can follow the formation of the VMS, as its mass is getting larger than the corresponding Lagrangian shell’s mass. Notably, the VMS reaches a mass of $M_{\rm VMS}\geq 1000\penalty 10000\ \rm M_{\odot}$ at time 0.180.18\,Myr (N-body simulation) and 0.220.22\,Myr (MOCCA simulation).
@@
Figure 3 shows the cumulative mass of the escapers during the simulation. In the first Myrs, the cumulative escaped mass is approximately $4\,000-5\,000\penalty 10000\ \rm M_{\odot}$, increasing to around $20\,000\penalty 10000\ \rm M_{\odot}$ after 55\,Myr. N-body and MOCCA simulation show reasonable agreement. The figure also presents the total number of collisions, which amounts to 22 01022\,010. Of these, $91.6\%$ involves the VMS/IMBH, these collisions can be divided into binary collisions ($45.1\%$) and hyperbolic collisions ($46.5\%$). Binary collisions are more frequent when the VMS mass exceeds $1\,000\penalty 10000\ \rm M_{\odot}$ at time 0.180.18\,Myr. In contrast, hyperbolic collisions become more frequent as the VMS grows in size, reaching a radius larger than $1\,000\penalty 10000\ \rm R_{\odot}$ at time 3.363.36\,Myr. For the MOCCA simulation, we do not distinguish between binary and hyperbolic collisions, as the code does not explicitly track stellar orbits. Some collisions may involve stars that were gravitationally bound to the VMS, but confirming this would require reconstructing stellar orbits from snapshots using their instantaneous positions, velocities, and the local gravitational potential, which is challenging due to the limited temporal resolution. Consequently, we report only the total number of collisions ($16103$ in 5 Myr), of which $65.6\%$ involve the VMS/IMBH. Despite this limitation, the results from both codes are in good agreement.
@@
Figure 4 compares the NBODY6++GPU and MOCCA simulations, showing the time evolution of the masses of escapers from the cluster (top panels) and secondary stars involved in collisions with the VMS/IMBH (bottom panels). The left panels display the N-body results, and the right panels show the MOCCA results. The color scale represents the local point density of events, with yellow regions indicating higher concentrations of recorded data points. In both simulations, low-mass stars (typically $\lesssim 1\penalty 10000\ \rm M_{\odot}$) dominate the population of escapers. The N-body simulation shows two prominent epochs of escaper activity: an extended early phase over the first $\sim 2\,$Myr, followed by a shorter but more intense burst near the time of IMBH formation at 4.694.69\,Myr. In contrast, the MOCCA simulation exhibits a smoother and more gradual decline in escaper activity over time, with a modest peak near $4.48$ Myr corresponding to IMBH formation.
@@
The lower panels show the masses of stars colliding with the VMS/IMBH. MOCCA displays a broader and more uniform spread across the stellar mass range from the start of the simulation, while the N-body simulation shows occasional high-mass collisions (with $m\gtrsim 10\penalty 10000\ \rm M_{\odot}$), though these are relatively sparse. In both simulations, most collisions involve low-mass stars. However, these frequent low-mass mergers contribute less significantly to the final VMS/IMBH mass (see also Figure 5). The N-body simulation shows a brief dip in the number of secondary stars involved in collisions around 1.681.68\,Myr, lasting approximately 0.10.1\,Myr. This feature coincides with the formation of a hard binary between the VMS and a main-sequence star (see Figure 7). Subsequent three-body interactions lead to enhanced escape rates. Toward the end of the simulation, the IMBH sequentially forms two hard binaries with $\sim 2\penalty 10000\ \rm M_{\odot}$ MS stars, which explains the observed drop in further collisions.
@@
Figure 5 shows histograms of the number of collisions (top panels) and the cumulative mass of the secondary star colliding with the VMS/IMBH (bottom panels). The N-body results are shown in the left panels, while the MOCCA results are shown in the right panels, the histogram mass ranges <0.1<0.1, 0.110.1-1, 1101-10, 1010010-100 and $100\penalty 10000\ \rm M_{\odot}$. The number of collisions for N-body are $2057$, $14525$, $2525$, $923$, $95$, while for MOCCA are $1053$, $7797$, $2273$, $1570$, $74$, for the first two mass ranges, N-body has about twice as many collisions as MOCCA, for the mass range $1-10\penalty 10000\ \rm M_{\odot}$ the number of collisions is about 300 more in N-body, while for the next mass range $10-100\penalty 10000\ \rm M_{\odot}$ it is higher for MOCCA, about 600 more collisions. Finally, for the last mass range $100\penalty 10000\ \rm M_{\odot}$ the number of collisions is slightly large in the N-body simulation. In the case of N-body simulations where it is possible to distinguish between binary (red bars) and hyperbolic (blue bars) collisions, we can observe that hyperbolic collisions contribute slightly more to the VMS formation for the mass range <0.1<0.1, $0.1-1\penalty 10000\ \rm M_{\odot}$. The mass range $1-10\penalty 10000\ \rm M_{\odot}$ shows almost the same number of collisions, however, binary collisions contribute slightly more to the VMS mass, while binary collisions contribute quite a bit more to the VMS formation for the mass range 1010010-100, and $100\penalty 10000\ \rm M_{\odot}$. Both codes shows the largest number of collisions comes from stars in the mass range $0.1-1\penalty 10000\ \rm M_{\odot}$. However, the largest mass contribution to the VMS comes from stars in the mass range $10-100\penalty 10000\ \rm M_{\odot}$, for MOCCA the mass contribution is $4.41\times 10^{4}\penalty 10000\ \rm M_{\odot}$, while in N-body it is $3.35\times 10^{4}\penalty 10000\ \rm M_{\odot}$ ($2.73\times 10^{4}\penalty 10000\ \rm M_{\odot}$ from binary collisions (cyan bar) and $6.34\times 10^{3}\penalty 10000\ \rm M_{\odot}$ from hyperbolic collisions (magenta bar)), the second large mass contribution to the VMS comes from the mass range $100\penalty 10000\ \rm M_{\odot}$, in MOCCA with $9.46\times 10^{3}\penalty 10000\ \rm M_{\odot}$ and in N-body with $1.26\times 10^{4}\penalty 10000\ \rm M_{\odot}$, with $1.17\times 10^{4}$ and $914.19\penalty 10000\ \rm M_{\odot}$, from binary and hyperbolic collisions, respectively. For the lower mass ranges <0.1<0.1, 0.110.1-1 and, $1-10\penalty 10000\ \rm M_{\odot}$ the VMS mass contribution in the same order is $94.49$, $2.59\times 10^{3}\penalty 10000\ \rm M_{\odot}$ and, $8.18\times 10^{3}\penalty 10000\ \rm M_{\odot}$ for MOCCA, while for N-body it is $184.01$, $4.69\times 10^{3}\penalty 10000\ \rm M_{\odot}$ and, $6.80\times 10^{3}\penalty 10000\ \rm M_{\odot}$, for this mass ranges the VMS mass contribution from binary and hyperbolic collisions is quite similar.
Figure 5: Histograms showing the contribution of colliding stars to VMS formation, divided into binary, hyperbolic, and total collisions. The mass ranges displayed are: $\penalty 10000\ \mathrm{M}_{\odot}, 0.10.1$$\penalty 10000\ \mathrm{M}_{\odot}, 11$$\penalty 10000\ \mathrm{M}_{\odot}, 1010$$100\penalty 10000\ \mathrm{M}_{\odot}$, and $100\penalty 10000\ \mathrm{M}_{\odot}$. Results are shown for both the N-body simulation (left column) and the MOCCA simulation (right column). The top panel displays the number of collisions, while the bottom panel shows the cumulative mass contributed by these collisions.
@@ 4.2 N-body results: VMS evolution and IMBH formation
In our N-body simulation, several collisions along the simulation to build up a VMS with a mass of $56996.91\penalty 10000\ \rm M_{\odot}$, eventually collapsing and forming an IMBH with a mass of $51625.99\penalty 10000\ \rm M_{\odot}$ at time 4.694.69\,Myr.
@@
Figure 6 shows the scatter of the position relative to the star cluster’s density center of the VMS (and IMBH thereafter), $\xi_{\mathrm{VMS,\,IMBH}}\penalty 10000\ \rm[pc]$, and the speed of the VMS (and IMBH, when formed), $\nu_{\mathrm{VMS,\,IMBH}}\penalty 10000\ [\mathrm{km\,s}^{-1}]$. Initially, the star with a mass of $130.58\penalty 10000\ \rm M_{\odot}$, which later becomes the VMS, is located at $0.088\rm pc$ from the center, sinking with a velocity of $32.9\penalty 10000\ \rm{km\,s}^{-1}$. After around 100 years, it moves closer to the center at $0.061\rm pc$, while its velocity increases nearly threefold to $99.7\penalty 10000\ \rm{km\,s}^{-1}$. A few hundreds years later, it reaches a velocity of $165.1\penalty 10000\ \rm{km\,s}^{-1}$ at a position of $0.018\rm pc$. This rapid infall toward the center is driven by the high cluster density, however, the VMS position and velocity oscillations are consistent with the orbital time. After two collisions with MS stars (i.e., stars with a fraction of a solar mass), the first significant collision occurs with a star of mass $54.89\penalty 10000\ \rm M_{\odot}$ following a binary interaction. At this point, the star is at $0.006\rm pc$. As shown in the figure, the motion of the VMS, in terms of position and velocity, first aligns with the expectations of mass segregation (steady decrease in distance to the center and a velocity profile that initially rises before declining) and later transitions to Brownian motion. At 1.681.68\,Myr a hard binary system forms between the VMS and an MS star. Figure 7 shows the semi-mayor axis and eccentricity of the hard binary formed between the VMS with a mass of $28\,891\penalty 10000\ \rm M_{\odot}$ and a secondary star with a mass of $75\penalty 10000\ \rm M_{\odot}$. Both stars are tightly bound due to their small average distance (<3<3 AU), they have an elliptical orbit, this dynamical interaction lasted about 0.10.1\,Myr. The binary system have a higher binding energy than the kinetic energy of the surrounding stars, which eventually gain kinetic energy, increasing their velocity. The latter reduces the probability of low-velocity stars that could collide with the VMS. Throughout the simulation, the velocity and position of the VMS remain nearly constant, fluctuating around an average value of the order of a $10^{-3}\rm pc$, while the velocity remains in moderate motion, at a few kilometers per second.
@@
Figure 7: The Figure shows the temporal evolution of a hard binary system during between 1.68 and 1.80 Myr. This binary has the primary star as the VMS, with a mass of $28891\penalty 10000\ \rm M_{\odot}$ and a secondary star with a mass of $75\penalty 10000\ \rm M_{\odot}$. This interaction lasts approximately $0.1$ Myr. The top panel shows the semi-major axis, while the bottom panel shows the eccentricity.
@@
As supported by the theoretical expectations discussed in previous sections, the VMS growth in the N-body simulation occurs on a short time-scale. The VMS original progenitor passes from an initial mass of $130.58\penalty 10000\ \rm M_{\odot}$ to exceed $10^{3}\penalty 10000\ \rm M_{\odot}$ after only 0.180.18\,Myr through 17 collisions, further reading $5000\penalty 10000\ \rm M_{\odot}$ in a time 0.290.29\,Myr. The VMS continues its growth for 0.480.48\,Myr through several collisions until reaching more than $10^{4}\penalty 10000\ \rm M_{\odot}$, naturally, the size of the VMS also increases having $71.13\rm R_{\odot}$, the continued collisions allow a strong rejuvenation of the star which has an effective age 0.230.23\,Myr. Lately, around 1.01.0\,Myr the VMS has a mass, effective age, and radius of $1.9\times 10^{4}\penalty 10000\ \rm M_{\odot}$, 0.540.54\,Myr, and $80.75\rm R_{\odot}$, respectively. The evolution of the mass, radius, and age of the VMS is shown in the top, middle, and bottom left panels of Figure 8.
@@
Figure 8: Evolution of the VMS (and IMBH thereafter) over time. Top panel: Mass of the VMS/IMBH, $M_{\mathrm{VMS\,IMBH}}\penalty 10000\ [\mathrm{M}_{\odot}]$. Middle panel: Stellar radius of the VMS/IMBH, $R_{\mathrm{VMS\,IMBH}}\penalty 10000\ [\mathrm{R}_{\odot}]$ (logarithmic scale). Bottom panel: Effective age of the VMS during its main-sequence phase, $A_{\mathrm{VMS\,IMBH}}\penalty 10000\ [\mathrm{Myr}]$.
@@
At 22\,Myr, the VMS has a mass of $3.1\times 10^{4}\penalty 10000\ \rm M_{\odot}$, with a radius $146.24\rm R_{\odot}$ and effective age 1.211.21\,Myr, the hard binary system formed at 1.681.68\,Myr in the N-body simulation prevents the VMS from experiencing further collisions, thus leads to a VMS with about $4\,000\penalty 10000\ \rm M_{\odot}$ less mass than the VMS obtained in MOCCA simulation, that has a mass of $3.5\times 10^{4}\penalty 10000\ \rm M_{\odot}$ at the same simulation time, 11\,Myr later the N-body VMS reaches a mass, radius, and effective age of $4.3\times 10^{4}\penalty 10000\ \rm M_{\odot}$, $507.24\rm R_{\odot}$ and 1.771.77\,Myr, respectively. The N-body VMS still have less mass than the MOCCA VMS which have a mass $4.52\times 10^{4}\penalty 10000\ \rm M_{\odot}$ at 33\,Myr.